# Low-metallicity massive single stars with rotation

### v2

@@ 5.2 Caveats
Caveats in our theory include uncertainties inherent in the simulations (see Sect. 6.2 of Paper II for a detailed discussion on this, as well as Agrawal et al. 2022). Some of the main sources of uncertainty are mass loss, wind clumping and rotational mixing efficiency. Especially mass loss and wind clumping can directly influence the strength of the spectral lines—but not the ionizing flux. For example, reducing the mass-loss rate by two orders of magnitude erases any trace of the carbon emission even in the post-main-sequence phase of our models, as seen in Fig. B4 of Paper II, however, this has practically no effect on the total number of ionizing photons, cf. Fig. B1 of the same paper. On the other hand, supposing that winds are clumped has an opposite effect on the emission lines: making them more prominent, as shown by Figs. 6 and 15-16 while not influencing the SED. This means that if mass-loss rates turn out to be lower then the rather high value assumed here, increasing wind clumping—another unconstrained property, cf. Roy et al. (2025)—may still lead to the same conclusion, making our results quite robust.
@@
As for the efficiency of rotational mixing, it is a widely studied yet unconstrained parameter in stellar evolution theory (Maeder & Meynet 2000; Meynet & Maeder 2000; Brott et al. 2011; de Mink et al. 2013; Burssens et al. 2023). However, our results are rather robust in this context as well. As explained in Sect. 2.3, we make the simple assumption that 10% of stars rotate fast enough for chemically-homogeneous evolution (guided by the observed initial rotational-velocity distribution of massive single stars in the SMC). So even if rotational mixing is found to be less strong than in our models, our results would be unchanged as long as 10% of the population follows the chemically-homogeneous path.
@@
The same is true for other modelling ingredients that influence stellar evolution (e.g. mass-loss rate prescriptions, convective overshooting/semi-convection, core-envelope coupling, etc.): with 10% of massive stars in the population evolving chemically homogeneously, either as single stars or binaries666This ratio is in line with Dorozsmai et al. (2024) who investigated populations of hierarchical triples with a chemically-homogeneously evolving inner binary (based on the rapid binary population synthesis method developed by Riley et al. 2021): in their Table 1, they report 10% of their systems following the chemically homogeneous path. (Whether such triples survive until core collapse is another question, as shown by Vigna-Gómez et al. 2025.), their spectra will look like how the PoWR models predict: so they will help us explain the peculiar features of I Zw 18.
@@
As shown recently by the seminal work of Abdellaoui (2023, cf. their Sect. 6.1.3), fast rotation has a strong effect on stellar winds, as well as on the emergent ionizing fluxes. Including these effects (by recomputing our PoWR models using the mass-loss rates obtained by Abdellaoui 2023) will be part of a future work. To repeat, we do not have any “classically” occurring WO stars in our population (i.e. WOs which got stripped by stellar-wind mass loss) because the evolutionary models we rely on do not predict such strong stellar winds at this low metal content (Szécsi et al. 2015, 2022). All our WOs evolved via the chemically-homogeneous route, as explained in Sect. 2.3.
@@
The star-formation history of I Zw 18 is not conclusively established (Kunth & Östlin 2000; Papaderos et al. 2002; Izotov & Thuan 2004; Papaderos & Östlin 2012), and indeed it has been debated whether it is a typical dwarf galaxy at all (Aloisi et al. 1999; Guseva et al. 2000; Annibali et al. 2013). For example, the initial mass function may be top-heavy in starbursts; this would help our case. Still, our result that massive stars of I Zw 18-composition are consistently able to account for two, previously unreconcilable observational characteristics while fulfilling all observational constraints available in the literature, implies that their existence should be supposed—and, therefore, further studied—in other galaxies of this type too. For example, some dwarf galaxies may be different from I Zw 18 not because of the typical evolutionary behaviours of their massive stars, but because their star formation history progressed another way due to e.g. different dynamical interactions with other galaxies (cf. Christensen et al. 2016), or to e.g. different thermodynamic/magneto-hydrodynamic conditions in their interstellar gas (cf. Hopkins et al. 2011; Seifried et al. 2017; Girichidis et al. 2018; Haid et al. 2019), leading to different rotational velocities (or different binary fraction) in the current massive stellar population. Indeed, to obtain a comprehensive theory of low-metallicity massive stars, one will need to correct for such environmental characteristics when accounting for a larger sample of dwarf galaxy observations (cf. also the conclusions in Stanway & Eldridge 2019), including any possible contribution of the interstellar medium to the emission lines commonly attributed to stellar origin.
@@ 6 Conclusions
We have shown that a population of massive Pop-II stars can consistently reproduce the observationally derived He II ionizing flux and the observed C IV 1550 Å line luminosity in the dwarf galaxy I Zw 18. The source of the emission bump is a small number of chemically-homogeneously evolved WO stars: these, combined with the rest of the hot O star population (both from normal evolution and chemically-homogeneous evolution, the latter dubbed ‘TWUIN’ in Paper I and Paper II in order to raise attention to their predicted low wind-opacity and thus the lack of emission features) explain the high photoionization. The evolutionary and atmospheric models were created with the measured composition of the gas in I Zw 18, thereby making our results self-consistent. Additionally, we carefully evaluated the literature and showed that neither the existing UV spectra nor the optical spectra are able to contradict our scenario.
@@
When performing population and spectral synthesis, we assumed a normal Salpeter mass function up to 120 M⊙ and a rather small fraction (10%) of single stars rotating fast enough for chemically-homogeneous evolution. This makes our result quite robust, as it means that the same conclusions would be drawn by a binary-population study too, as long as 10% of the systems evolve chemically homogeneously. Varying assumptions about the star-formation history, the best match predicts twenty WO stars of mass around 25 M⊙ and two of 50 M⊙ to account for the observations, inlcuding the spectral hardness.
@@
Furthermore, we have established a new sequence of evolutionary phases, typical for chemically-homogeneously evolving stars at Z=0.0002: O WN WO. The WC phase, if arises at all between WN and WO, lasts no longer than 1% of the total lifetime. While we have based this study on single stars, we expect the same sequence for binaries in the chemically-homogeneously evolving channel (Hainich et al. 2018).
@@
From our careful review of the literature, we conclude that while certain kind of signal may dominate in one measurement (like how stellar emission dominates in some of the spectra we discuss in Sect. 4), it does not follow that it will dominate all other measurements from the same, spatially extended galaxy. The telescopes may point at different regions, apply differing extraction apertures, or have lower sensitivity in a given spectral range. Paying attention to the observational strategy (e.g. slit vs. integrated spectrum, pointing, size of aperture, etc.) is crucial when comparing various pieces of data from various instruments, especially when it comes to spatially resolved objects such as I Zw 18. In this context, detecting the emission bump around O VI 3818 Å (where the existing data is problematic, Sect. 4.1) is a worthwhile future goal that could conclusively prove our theory about these WO stars—and thus, the existence of chemically-homogeneous evolution. At the faintness of I Zw 18, such a measurement can only be done by a 6-10m class telescope with a low/medium resolution spectrograph.
@@
Our stars are expected to end their lives with certain explosions such as long-duration gamma-ray bursts (Yoon & Langer 2005; Woosley & Heger 2006; Yoon et al. 2006), supernovae of type I b/c or even superluminous supernovae of the hydrogen-poor type I (Szécsi 2016, 2017a, 2017b; Aguilera-Dena et al. 2018, which of these would happen, depends mainly on their mass). In a close binary system they may even form a double compact object which leads to gravitational wave emission upon merging (de Mink et al. 2009; de Mink & Mandel 2016; Mandel & de Mink 2016; du Buisson et al. 2020; Riley et al. 2021; Sharpe et al. 2024). Finding evidence for the existence of this evolutionary pathway in I Zw 18 is, therefore, of upmost importance, and serves as motivation to study this dwarf galaxy further in the future.
@@
Our research on metal-poor massive stars in I Zw 18 has implications not only for local dwarf galaxies, but for other types of galaxies as well. For example, it has been suggested (Micheva et al. 2017) that the so-called Green Pea galaxies (Izotov et al. 2011; Jaskot & Oey 2013; Yang et al. 2016; Orlitová et al. 2018; Franeck et al. 2022; Arroyo-Polonio et al. 2023; Smith et al. 2023) are dwarf-galaxy analogues at intermediate cosmological distances; if so, their properties may only be understood by including chemically-homogeneously evolving massive stars into the picture. Similarly, our knowledge of high-redshift galaxies e.g. in the epoch of cosmic reionization may need to be re-evaluated in light of what chemically-homogeneous stars can bring onto the table in terms of ionizing radiation (Eldridge & Stanway 2012; Sobral et al. 2015; Stanway et al. 2016; Visbal et al. 2016; Sobral et al. 2019) and other types of stellar feedback (Bowler et al. 2017). Given the recent result of Liu et al. (2025) showing—with a completely independent method from ours—that chemically-homogeneously evolving stars may indeed exist in early cosmic epochs, our conclusions have far-reaching implications for future surveys with JWST and other campaigns targeting the dawn of our Universe.

### v3

@@ 5.2 Caveats
Caveats in our theory include uncertainties inherent in the simulations (see Sect. 6.2 of Kubatova:2019 for a detailed discussion on this, as well as Agrawal:2022). Some of the main sources of uncertainty are mass loss, wind clumping and rotational mixing efficiency. Especially mass loss and wind clumping can directly influence the strength of the spectral lines – but not the ionizing flux. For example, reducing the mass-loss rate by two orders of magnitude erases any trace of the carbon emission even in the post-main-sequence phase of our models, as is seen in Fig. B4 of Kubatova:2019; however, this has practically no effect on the total number of ionizing photons (see Fig. B1 of the same paper). On the other hand, supposing that winds are clumped has an opposite effect on the emission lines: making them more prominent, as is shown by Figs. 6 and 15-16 while not influencing the SED. This means that if mass-loss rates turn out to be lower then the rather high value assumed here, increasing wind clumping – another unconstrained property (see Roy:2025) – may still lead to the same conclusion, making our results quite robust.
@@
As for the efficiency of rotational mixing, it is a widely studied yet unconstrained parameter in stellar evolution theory (Maeder:2000; Meynet:2000; Brott:2011a; deMink:2013; Burssens:2023). However, our results are rather robust in this context as well. As is explained in Sect. 2.3, we make the simple assumption that 10% of stars rotate fast enough for CHE (guided by the observed initial rotational-velocity distribution of massive single stars in the SMC). So even if rotational mixing is found to be weaker than in our models, our results would be unchanged as long as 10% of the population follows the chemically homogeneous path.
@@
The same is true for other modelling ingredients that influence stellar evolution (e.g. mass-loss rate prescriptions, convective overshooting, semi-convection, core-envelope coupling): with 10% of massive stars in the population evolving chemically homogeneously, either as single stars or binaries777This ratio is in line with Dorozsmai:2024 who investigated populations of hierarchical triples with a chemically homogeneously evolving inner binary (based on the rapid binary population synthesis method developed by Riley:2021): in their Table 1, they report 10% of their systems following the chemically homogeneous path. (Whether such triples survive until core collapse is another question, as is shown by Vigna-Gomez:2025.), their spectra will look like how the PoWR models predict: so they will help us explain the peculiar features of I Zw 18.
@@
As was shown recently by the seminal work of Abdellaoui:2023, fast rotation has a strong effect on stellar winds, as well as on the emergent ionizing fluxes. Including these effects (by recomputing our PoWR models using the mass-loss rates obtained by Abdellaoui:2023) will be part of a future work. To repeat, we do not have any “classically” occurring WO stars in our population (i.e. WOs that got stripped by stellar-wind mass loss) because the evolutionary models we rely on do not predict such strong stellar winds at this low metal content (Szecsi:2015; Szecsi:2022ok). All our WOs evolved via the chemically homogeneous route, as explained in Sect. 2.3.
@@
The star formation history of I Zw 18 is not conclusively established (Kunth:2000; Papaderos:2002; Izotov:2004; Papaderos:2012), and indeed it has been debated whether it is a typical dwarf galaxy at all (Aloisi:1999; Guseva:2000; Annibali:2013). For example, the IMF may be top-heavy in starbursts; this would help our case. Still, our result that massive stars of I Zw 18-composition are consistently able to account for two, previously unreconcilable observational characteristics while fulfilling all observational constraints available in the literature, implies that their existence should be supposed – and, therefore, further studied – in other galaxies of this type too. For example, some dwarf galaxies may be different from I Zw 18 not because of the typical evolutionary behaviours of their massive stars, but because their star formation history progressed another way due to, for example, different dynamical interactions with other galaxies (see Christensen:2016), or to different thermodynamic/magneto-hydrodynamic conditions in their interstellar gas (see Hopkins:2011; Seifried:2017; Girichidis:2018; Haid:2019), leading to different rotational velocities (or different binary fraction) in the current massive stellar population. Indeed, to obtain a comprehensive theory of low-metallicity massive stars, one will need to correct for such environmental characteristics when accounting for a larger sample of dwarf galaxy observations (see also the conclusions in Stanway:2019), including any possible contribution of the interstellar medium to the emission lines commonly attributed to stellar origin.
@@ 6 Conclusions
We have shown that a population of massive Pop-II stars can consistently reproduce the observationally derived He II ionizing flux and the observed C IV 1550 Å line luminosity in the dwarf galaxy I Zw 18. The source of the emission bump is a small number of chemically homogeneously evolved WO stars: these, combined with the rest of the hot O star population (both from normal evolution and CHE, the latter dubbed ‘TWUIN’ in Szecsi:2015 and Kubatova:2019 in order to raise attention to their predicted low wind-opacity and thus the lack of emission features) explain the high photoionization. The evolutionary and atmospheric models were created with the measured composition of the gas in I Zw 18, thereby making our results self-consistent. Additionally, we carefully evaluated the literature and showed that neither the existing UV spectra nor the optical spectra are able to contradict our scenario.
@@
When performing population and spectral synthesis, we assumed a normal Salpeter mass function up to 120 M⊙ and a rather small fraction (10%) of single stars rotating fast enough for CHE. This makes our result quite robust, as it means that the same conclusions would be drawn by a binary-population study too, as long as 10% of the systems evolve chemically homogeneously. By varying our assumptions about the star formation history, the best match we found predicts 20 WO stars that have a mass of around 25 M⊙ and two of 50 M⊙ to account for the observations, including the spectral hardness.
@@
Furthermore, we have established a new sequence of evolutionary phases, typical for chemically homogeneously evolving stars at Z=0.0002: O WN WO. The WC phase, if it arises at all between WN and WO, lasts no longer than 1% of the total lifetime. While we have based this study on single stars, we expect the same sequence for binaries in the chemically homogeneously evolving channel (Hainich:2018).
@@
From our careful review of the literature, we conclude that while a certain kind of signal may dominate in one measurement (similarly to how stellar emission dominates in some of the spectra we discuss in Sect. 4), it does not follow that it will dominate all other measurements from the same, spatially extended galaxy. The telescopes may point at different regions, apply differing extraction apertures, or have lower sensitivity in a given spectral range. Paying attention to the observational strategy (e.g. slit vs integrated spectrum, pointing, size of aperture) is crucial when comparing various pieces of data from various instruments, especially when it comes to spatially resolved objects such as I Zw 18. In this context, detecting the emission bump around O VI 3818 Å (where the existing data is problematic, Sect. 4.1) is a worthwhile future goal that could conclusively prove our theory about these WO stars – and thus, the existence of CHE. At the faintness of I Zw 18, such a measurement can only be done by a 6-10m class telescope with a low- or medium-resolution spectrograph.
@@
Our stars are expected to end their lives with certain explosions such as long-duration gamma-ray bursts (Yoon:2005; Woosley:2006; Yoon:2006), supernovae of type I b/c, or even superluminous supernovae of the hydrogen-poor type I (Szecsi:2016; Szecsi:2017short; Szecsi:2017long; AguileraDena:2018, which of these would happen depends mainly on their mass). In a close binary system, they may even form a double compact object that leads to gravitational wave emission upon merging (deMink:2009b; deMink:2016; Mandel:2016; duBuisson:2020G; Riley:2021; Sharpe:2024G). Finding evidence of the existence of this evolutionary pathway in I Zw 18 is therefore of upmost importance, and serves as motivation to study this dwarf galaxy further in the future.
@@
Our research on metal-poor massive stars in I Zw 18 has implications not only for local dwarf galaxies but for other types of galaxies as well. For example, it has been suggested (Micheva:2017) that the so-called Green Pea galaxies (Izotov:2011; Jaskot:2013; Yang:2016; Orlitova:2018; Franeck:2022; ArroyoPolonio:2023; Smith:2023) are dwarf-galaxy analogues at intermediate cosmological distances; if so, their properties may only be understood by including chemically homogeneously evolving massive stars into the picture. Similarly, our knowledge of high-redshift galaxies, for example in the epoch of cosmic reionization, may need to be re-evaluated in light of what chemically homogeneous stars can bring to the table in terms of ionizing radiation (Eldridge:2012; Sobral:2015; Stanway:2016; Visbal:2016; Sobral:2019) and other types of stellar feedback (Bowler:2017). Given the recent result of Liu:2025 showing – with a completely independent method from ours – that chemically homogeneously evolving stars may indeed exist in early cosmic epochs, our conclusions have far-reaching implications for future surveys with JWST and other campaigns targeting the dawn of our Universe.